OpenStudy (anonymous):

cos x/2 = cos π. find all solutions in the interval [0,2π)?

OpenStudy (anonymous):

\[cos(\pi)=-1\] so your equation is \[cos(\frac{x}{2})=-1\] one solution is \[\frac{x}{2}=\pi\] so \[x=2\pi\] that is the only one in your interval. in fact it is not in your interval, so perhaps there is no solution!